The Carla Vista Acres Inn is trying to determine its break-even point during its off-peak season. The inn has 50 rooms that it rents at $70 a night. Operating costs are as follows: Salaries $6,000 per month Utilities $1,400 per month Depreciation $1,300 per month Maintenance $3,228 per month Maid service $14 per room Other costs $28 per room

Part 1) Determine the inn’s break-even point in number of rented rooms per month.

Part 2) Determine the inn’s break-even point in dollars.
Answer #1

Fixed cost = Salaries + Utilities + Depreciation + Maintenance

= 6,000+1,400+1,300+3,228

= $11,928

Variable cost per room = Maid service + Other costs

= 14+28

= $42

Room rent per night = $70

Contribution room per night = Room rent per night-Variable cost per room

= 70-42

= $28

Part 1

Break even point in number of rented rooms per month = Fixed cost / Contribution room per night

= 11,928/28

= 426 rented rooms per month

Part 2

Contribution margin ratio = Contribution room per night / Room rent per night

= 28/70

= 40%

Break-even point in dollars = Fixed cost / Contribution margin ratio

= 11,928/40%

=$29,820

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.

Thanks!!!